Crossfire Company segments its business into two regions—East and West. The company prepared a contribution format segmented income statement as shown below:

[The following information applies to the questions displayed below.]

  

Raner, Harris & Chan is a consulting firm that specializes in information systems for medical and dental clinics. The firm has two offices—one in Chicago and one in Minneapolis. The firm classifies the direct costs of consulting jobs as variable costs. A contribution format segmented income statement for the company’s most recent year is given:

Office
Total Company Chicago Minneapolis
Sales $ 900,000 100.0 % $ 180,000 100 % $ 720,000 100 %
Variable expenses 486,000 54.0 % 54,000 30 % 432,000 60 %
Contribution margin 414,000 46.0 % 126,000 70 % 288,000 40 %
Traceable fixed expenses 201,600 22.4 % 93,600 52 % 108,000 15 %
Office segment margin 212,400 23.6 % $ 32,400 18 % $ 180,000 25 %
Common fixed expenses not traceable to offices 144,000 16.0 %
Net operating income $ 68,400 7.6 %

Required:

1-a. Compute the companywide break-even point in dollar sales.

1-b. Compute the break-even point for the Chicago office and for the Minneapolis office.

1-c. Is the companywide break-even point greater than, less than, or equal to the sum of the Chicago and Minneapolis break-even points?
Answer #1

1-a Company wide breakeven point = ( Total traceable fixed expenses + Common fixed expenses ) / Contribution margin ratio of the company = ( $201,600 + $144,000 ) / 46% = $751,304.35

1-b Breakeven point for the Chicago office = Traceable fixed expenses of Chicago office / Contribution margin ratio of Chicago office = $93,600 / 70% = $133,714.29

Breakeven point for the Minneapolis office = Traceable fixed expenses of Minneapolis / Contribution margin ratio of Minneapolis = $108,000 / 40% = $270,000

1-c Sum of the Chicago and Minneapolis breakeven point = $133,714.29 + $270,000 = $403,714.29

Companywide breakeven point is ( $751,304.35 ) is greater than the sum of the Chicago and Minneapolis breakeven point ( $403,714.29 )